OpenStudy (anonymous):

please help How many solutions does it have? -3x+7y=8/3x+5y Does it have 2 or more than two?

OpenStudy (anonymous):

Its -3x+ 7y = 8 / 3x+ 5y=6

OpenStudy (anonymous):

Solve for y

OpenStudy (anonymous):

y=14/12

OpenStudy (anonymous):

if it can be solved for x and y is both two solutions?

OpenStudy (owlcoffee):

So these are are system of equations?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I wanna know if its infitite solutions or just 2

OpenStudy (owlcoffee):

So, let's call them equation 1 and equation 2 respectively, so, we'll have this: \[[1] -3x+7y=8\] \[[2] 3x+5y=6\] Now, there are many ways to solve this type of linear system but I choose to use the method of reduction, which means that I will operate these two equations in such a way that I will leave only one variable, making it easier to solve for. I will do that by adding them: \[[1]+[2] \iff (-3x+7y)+(3x+5y)=14\] \[12y=14\] \[y=\frac{ 7 }{ 6 }\] Look, the fact that we got a value for "y" implies that there is a fixed point as a solution, this means that it only has one solution.

OpenStudy (anonymous):

I just want to know one thng I kinda know the answer

OpenStudy (anonymous):

what if we multiply by 5 the first column wouldnt it give like x= 1/3?

OpenStudy (owlcoffee):

Show me what you mean.

OpenStudy (anonymous):

doesnt it give like -15 x + 35= 40 and then multiply for the second for -5 like -15x- 35 = -30 -30x=10 x=-10/30 x=-1/3

OpenStudy (owlcoffee):

I see, yes, that works as well.